OpenStudy (anonymous):

EASY MEDAL !!!!! Given the trinomial, what is the value of the coefficient 'B' in the factored form? 2x2 - 12xy - 32y2 = 2(x - 8y)(x + By) B= ____________ Numerical Answers Expected!

OpenStudy (loser66):

any idea?

OpenStudy (anonymous):

2(x^2-6xy-16y^2)=2(x-8y)(x+2y)

OpenStudy (anonymous):

is that the answer @Plee

OpenStudy (anonymous):

@Loser66 help

OpenStudy (loser66):

what do you want from me? answer? ok, it's B = 2 If you wonder how can I get it, let me know. If you are not, FINE

OpenStudy (anonymous):

how did you get it ?

OpenStudy (loser66):

I take it as you want to study, am I right?

OpenStudy (anonymous):

yeah

OpenStudy (loser66):

good, expand the right hand side, what do you get?

OpenStudy (anonymous):

2(x^2-6xy-16y^2)=2(x-8y)(x+2y)

OpenStudy (loser66):

nope, I talk about the original one, open the bracket of the right hand side

OpenStudy (anonymous):

Factor each trinomial. Then match the polynomial (term) on the left with its factored form (definition) on the right. Match Term Definition x2 – 8x – 20 A) (x – 5)(x + 4) x2 + 8x – 20 B) (x – 10)(x + 2) x2 – x – 20 C) (x – 2)(x + 10) x2 – 9x – 20 D) Prime

OpenStudy (anonymous):

@Loser66 can you help me on this one i will give you a medal

OpenStudy (loser66):

I don't work for medal, sorry, friend, feel free to take it off.

OpenStudy (loser66):

1 B 2c 1A 4 D enough for you, no more

OpenStudy (anonymous):

THANK YOU SO MUCH MAN

OpenStudy (anonymous):

i just did this for hmwk. the answer is 2